OpenStudy (anonymous):

I'm not sure if this one is -1/2 Solve for the variable 5z+4=z+(16/5)

OpenStudy (anonymous):

check it... (5.-1/2)+4=-1/2+(16/5) -10+8=-1+32/5 oh...check it urself...!

OpenStudy (anonymous):

Obviously I did something wrong, I need help finding z